# The impact of stellar bars on quenching star formation: insights from a spatially resolved analysis in the local Universe

### v1

@@ 2.3 Sample description
This paragraph focuses on outlining the main characteristics of the selected galaxies, building on previous literature findings and highlighting the most relevant to our analysis.
@@
Firstly, it is important to note that although we decided to rule out any possible strong AGN contamination, our sample includes two galaxies, namely NGC 1566 and NGC 4579, in which there are hints of activity in their nuclear regions. NGC 1566 is considered a low-luminosity AGN (LLAGN), but it shows a soft X-ray spectrum that is likely explained with a luminous nuclear starburst (Tomás et al., 2022). Thanks to ALMA observations, Combes et al. (2014) found that the HCO+(4-3) line is three times stronger than the HCN(4-3), as expected when star formation excitation dominates over AGN heating. Also, Vermot et al. (2023) suggested that the nucleus of this galaxy hosts both phenomena: very recent or ongoing star formation and an AGN (accretion disc + hot dust). The latter contribution is supported by detection in optical and NIR spectra of associated emission lines (Smajić et al., 2015; da Silva et al., 2017). This galaxy is characterised by variable nuclear activity and observations of the AGN in the outburst phase seem to identify disc instabilities as the cause of such variability (Parker et al., 2019). NGC 4579 is classified as an intermediate LINER/Seyfert 1.9 spiral galaxy, according to the classification of Ho et al. (1997). Subsequent works, e.g., García-Burillo et al. (2009), considered this object a representative example of a LLAGN in which several mechanisms can co-operate to transport the gas from the outer disc to the central regions (spiral arms, bar, viscosity), thus feeding star formation and possibly the AGN activity itself. More recently, this galaxy has been studied in terms of variability timescales: Younes et al. (2019) found a lack of variability in the X-ray emission and a broad Fe K line. They argued for an altered accretion geometry in LLAGN, compared to luminous AGN.
@@
We decided to include both of these objects in our sample because their LLAGN appears to have only a marginal impact on the overall galaxy properties and on its energy budget (Casasola et al., 2015). This nuclear activity could at most affect the inner sub-kpc region. For example, de Sá-Freitas et al. (2023) masked a region of at the centre of NGC 1566 in order to avoid any potential contamination in their analysis, which was based on the derivation of kinematic and stellar population properties in the nuclear disc. This angular scale is smaller than our typical aperture employed in the subsequent spatially resolved procedure (). For this reason, we initially assume that AGN activity in their nuclear region does not affect our data analysis, and then we verify this assumption a posteriori by looking at the SED fitting in the central pixels. (see Sect. 5).
@@
NGC 3351 is actively forming stars in its nucleus and in a circumnuclear ring (with a diameter of 0.7 kpc), as it is highlighted by many multi-wavelength studies performed on sub-kpc scales (Elmegreen et al., 1997; Bresolin & Kennicutt, 2002; Ma et al., 2018). H imaging of this galaxy shows that the bar region is devoid of emission (James et al., 2009) and, according to James & Percival (2016), it is characterised by an old stellar population. So far, literature has suggested that the nuclear starburst and the observed suppression of recent star formation () in the bar region, which lacks neutral and molecular hydrogen, are likely due to the influence of the bar itself (e.g., George et al., 2019a).
@@
On the contrary, there are not many studies concerning NGC 3953 and NGC 4725. The former has a rather low bulge/disc ratio, suggesting that its inner regions are mostly dominated by components other than the bulge, such as the bar and ring (Castro-Rodríguez & Garzón, 2003). Besides, the work of Staudaher et al. (2019) seems to show the presence of some leftover turbulent gas in this galaxy, resulting in a newly discovered looping tidal stream. The latter, i.e., NGC 4725, was identified as a double-barred galaxy by Erwin (2004), but de Lorenzo-Cáceres et al. (2013) suggested that, at present, the inner bar is playing a minor role in promoting major morphological changes. Chiaraluce et al. (2019) attempted to measure its nuclear radio emission, including this object among a sample of low-luminosity Seyfert galaxies (Véron-Cetty & Véron, 2006). However, their Very Large Array (VLA) observations produced an unresolved image and a peculiar radio spectrum, which would deserve further investigation to derive meaningful information from this source.
@@
Regarding NGC 5236, it has a grand design barred spiral morphology and hosts a nuclear starburst ring (Buta & Crocker, 1993; Calzetti et al., 2004; Knapen et al., 2010) leading to a high total star formation rate (SFR 4.2 yr-1, Leroy et al., 2021). Furthermore, Della Bruna et al. (2022) observed that both stars and gas kinematics trace regular rotation along the main star-forming disc as well as a fast-rotating circumnuclear disc of ( 700 pc) in diameter, likely originating from secular processes driven by the galactic bar (as already reported by Gadotti et al. (2020)). They also detected an extended ionised gas region, which could, among other factors, be attributed to a bar-driven inflow of shocked gas.
@@ 2.4 Environmental properties
The environment in which galaxies inhabit is believed to play a crucial role in their evolutionary paths (Dressler, 1980; Poggianti et al., 1999). According to many observational works, late-type galaxies (spiral, blue and star forming) are more likely found in lower density environments, than early-type ones (elliptical, red and quenched), which are more numerous in groups, clusters and interacting systems (e.g., Casasola et al., 2004; Blanton et al., 2005; Peng et al., 2010; Boselli & Gavazzi, 2014). Indeed, there are several external processes that are commonly identified as responsible for this phenomenology and that can alter the effects of internal and/or secular processes, such as the action of a stellar bar. Among them, we can mention: ram-pressure stripping (Gunn & Gott, 1972; Yun et al., 2019), mergers (Mihos & Hernquist, 1994), harassment (high-speed encounters between galaxies; Moore et al., 1996), strangulation (cutoff of gas accretion; Kawata & Mulchaey, 2008). In this paragraph we present the environment in which each of our barred galaxies is located, in order to find out any possible influence.
@@
NGC 1566 is the brightest member of the Dorado group (Agüero et al., 2004; Kilborn et al., 2005) and belongs to a pair-like structure where its companion is the dwarf elliptical NGC 1581 Rampazzo et al. (2022). Moreover, the work of Elagali et al. (2019) found that NGC 1566 has an asymmetric and mildly warped HI disk. To explain this feature, they excluded the scenario of past interactions with other galaxies in favour of the possibility of having experienced ram-pressure stripping from the Intergalactic Medium (IGM), at least in its outskirts. However, the high atomic gas fraction that characterises this galaxy hints at a subtle ram-pressure stripping effect, as expected in galaxies’ groups.
@@
We used the catalogue produced by Karachentsev et al. (2013) to get the environmental properties of NGC 3351 and NGC 5236. This catalogue collects information about 869 galaxies within 11 Mpc and includes three estimates of the “tidal index”, which quantifies the local density environment. In particular, we looked at the tidal index , determined by the five most important neighbours. Positive values of this parameter indicate the membership in groups, while negative values correspond either to isolated galaxies or to galaxies in the outskirts of a group. NGC 3351 has a relatively high value of compared to the distribution of galaxies in the Karachentsev et al. (2013), meaning that it is in a mildly dense environment. On the other hand, NGC 5236 has and so it is a “field” galaxy. In support of this definition, we mention that NGC 5236 is present in the catalogue of isolated galaxies of Bettoni et al. (2003).
For what concerns NGC 3953, Kourkchi & Tully (2017) identified it as part of a group, whose main member is the barred spiral NGC 3992.
@@
According to Vollmer et al. (2012), NGC 4579 is one of the brightest galaxies in the Virgo Cluster. It is an anemic galaxy and has a relatively small amount of neutral hydrogen, located inside its disk and concentrated in clumps (Cayatte et al., 1990). This deficiency of gas content is thought to be a consequence of interactions with the Virgo Cluster’s intracluster medium, suggesting that ram-pressure stripping may have occurred in the galaxy’s outskirts (Koopmann & Kenney, 2004).
@@
NGC 4725 is the brightest member of the Coma I aggregation (Gregory & Thompson, 1977) and seems to be interacting with the neighbouring spiral galaxy NGC 4747 (Haynes, 1979). Gravitational interactions between these two galaxies likely resulted in a gas-rich tidal tail that extends about ( 30 kpc) to the north-east of NGC 4747 and appears to contain two distinguishable star clusters (Lee-Waddell et al., 2016, 2018). On the other hand, only a relatively small distortion in the HI distribution is observed in NGC 4725.
@@ 2.5 Multi-wavelength observations
As briefly aforementioned, DustPedia provides us with multi-wavelength observations, from ultraviolet to microwave. For both Spectral and Photometric Imaging REceiver (SPIRE; Griffin et al., 2010) and Photodetector Array Camera and Spectrometer (PACS; Poglitsch et al., 2010) data, the Herschel Science Archive (HSA555https://archives.esac.esa.int/hsa/whsa/) was queried to find all Herschel photometric observations covering each DustPedia galaxy. The ancillary imaging data were gathered from other seven facilities that observed large numbers of nearby galaxies: the GALaxy Evolution eXplorer (GALEX; Morrissey et al., 2007), the Sloan Digital Sky Survey (SDSS; York et al., 2000; Eisenstein et al., 2011), the Digitized Sky Survey (DSS), the 2 Micron All-Sky Survey (2MASS Skrutskie et al., 2006), the Wide-field Infrared Survey Explorer (WISE Wright et al., 2010), the Spitzer Space Telescope (Werner et al., 2004) and Planck (Planck Collaboration I, 2011). More information on the different data sets available can be found in Enia et al. (2020) and references therein.
Thanks to the data derived from these multi-wavelength observations collected in the DustPedia archive, it is possible to exploit the power of such a complete multi-band photometric coverage, especially in providing a measure of the comprehensive SFR in galaxies, directly taking into account both the optical and dust-obscured components. For what concerns our work, we use observations from GALEX, SDSS, 2MASS, WISE, Spitzer and Herschel, going from the FUV (1516 Å) to the FIR (350 m). However, it has to be noticed that not all of these data sets are accessible for the whole sample. Indeed, all 23 photometric bands are available only for NGC 3351, NGC 4579 and NGC 4725. NGC 3953 is observed in 20 bands; GALEX filters (FUV, NUV) and PACS 70m photometry are not available for this object. Instead, 18 bands are accessible for NGC 1566 and NGC 5236, which lack optical observations (the five Sloan optical filters).
In Fig. 1 we show the selected galaxies as they are observed in six different bands: FUV with GALEX, g-band with SDSS, J-band with 2MASS, 4.5 m with Spitzer, 12 m with the WISE and 250 m with Herschel. An elongated morphological structure is clearly visible in all the galaxies, both in the SDSS optical bands (when available) and in the NIR ones (2MASS and Spitzer).
@@ 3 Data analysis
The primary goal of our data analysis is to obtain the spatially resolved distributions of the two galaxy physical properties: stellar mass and star formation rate. In this respect, we carry out a SED fitting procedure on sub-kpc scales, making use of the publicly available magphys code (da Cunha et al., 2008). The methodology that we adopt to perform the data analysis has been developed and already used in previous works for a sample of local star forming galaxies with a “grand-design” spiral structure (Enia et al., 2020; Morselli et al., 2020). Our aim is to ensure a fair comparison with their results about unbarred galaxies, since they have consistent global properties with respect to ours.
In Enia et al. (2020) the procedure is summarised in three main steps: (i) degradation of the resolution of each image to the worst Point Spread Function (PSF) that is available among all the photometric bands (i.e., that of the m image taken with the SPIRE instrument, characterised by a PSF FWHM of ); (ii) set up of a grid of square cells of fixed size from which to measure the flux at different wavelengths; (iii) derivation of the physical properties for individual apertures by performing SED fitting to the available photometric data. We analyse every selected source considering cells of (corresponding to physical scales varying from to , as reported in Table 1), which is the pixel scale of SPIRE m maps. This choice is justified by the fact that we need a good balance between resolution and number of available photometric bands, in order to properly sample the SED, including the dust emission curve, and avoid a too low resolution that would not allow us to investigate the bar region. Specifically, it is necessary to sample the bar radius of each galaxy through, at least, three resolution elements (cells of each), to get accurate information about this morphological structure.
At this point, we highlight the fact that, throughout the paper, we refer to the “region hosting the bar” as the approximately annular region around the very central nucleus that extends up to the radius of the bar.
@@ 3.1 Pre-processing and PSF degradation
The preliminary steps that are performed on each photometric observation are those related to the pre-processing and PSF homogenisation phases, which start with the removal of foreground stars by masking contaminating sources through an automated procedure embedded in CAAPR. Subsequently, a background estimation and subtraction are performed on each of the star-subtracted maps following the guidelines provided by Clark et al. (2018). Finally, the resolution of all the stars and background subtracted maps is degraded to the PSF of SPIRE 350, exploiting the kernels of Aniano et al. (2011).

### v2

@@ 2.3 Sample description
Here, we outline the main characteristics of the selected galaxies, building on previous literature findings and highlighting those most relevant to our analysis. Firstly, it is important to note that although we decided to rule out any possible strong AGN contamination, our sample includes two galaxies in which there are hints of activity in their nuclear regions, namely NGC 1566 and NGC 4579. NGC 1566 is considered a low-luminosity AGN (LLAGN), but it shows a soft X-ray spectrum that is likely explained with a luminous nuclear starburst (Tomás et al., 2022). Thanks to ALMA observations, Combes et al. (2014) found that the HCO+(4-3) line is three times stronger than the HCN(4-3), as expected when star formation excitation dominates over AGN heating. Also, Vermot et al. (2023) suggested that the nucleus of this galaxy hosts both phenomena: very recent or ongoing star formation and an AGN (accretion disc + hot dust). The latter contribution is supported by detection in optical and NIR spectra of associated emission lines (Smajić et al., 2015; da Silva et al., 2017). This galaxy is characterised by variable nuclear activity and observations of the AGN in the outburst phase seem to identify disc instabilities as the cause of the variability (Parker et al., 2019). NGC 4579 was classified as an intermediate LINER/Seyfert 1.9 spiral galaxy by Ho et al. (1997). Subsequent works, such as García-Burillo et al. (2009), considered this object a representative example of a LLAGN, in which several mechanisms can co-operate to transport the gas from the outer disc to the central regions (spiral arms, bar, viscosity), thus feeding star formation and possibly the AGN activity itself. More recently, this galaxy was studied in terms of variability timescales: Younes et al. (2019) found a lack of variability in the X-ray emission and a broad Fe K line, arguing for an altered accretion geometry in LLAGN compared to luminous AGN.
@@
We decided to include both of these objects in our sample because their LLAGN appears to have only a marginal impact on the overall galaxy properties and on its energy budget (Casasola et al., 2015). This nuclear activity could at most affect the inner subkpc region. For example, de Sá-Freitas et al. (2023) masked a region of at the centre of NGC 1566 in order to avoid any potential contamination in their analysis, which was based on the derivation of kinematic and stellar population properties in the nuclear disc. This angular scale is smaller than our typical aperture employed in the subsequent spatially resolved procedure (). For this reason, we initially assumed that AGN activity in their nuclear region does not affect our data analysis, and then we verified this assumption a posteriori by looking at the SED fitting in the central pixels (see Sect. 5).
@@
NGC 3351 is actively forming stars in its nucleus and in a circumnuclear ring (with a diameter of 0.7 kpc), as is highlighted by many multi-wavelength studies performed on subkpc scales (Elmegreen et al., 1997; Bresolin & Kennicutt, 2002; Ma et al., 2018). H imaging of this galaxy shows that the bar region is devoid of emission (James et al., 2009) and, according to James & Percival (2016), is characterised by an old stellar population. The existing literature suggests that the nuclear starburst and the observed suppression of recent star formation () in the bar region —which lacks neutral and molecular hydrogen— are likely due to the influence of the bar itself (e.g. George et al., 2019a).
@@
On the contrary, there are very few studies concerning NGC 3953 and NGC 4725. The former has a rather low bulge/disc ratio, suggesting that its inner regions are mostly dominated by components other than the bulge, such as the bar and ring (Castro-Rodríguez & Garzón, 2003). The work of Staudaher et al. (2019) seems to show the presence of some leftover turbulent gas in this galaxy, resulting in a newly discovered looping tidal stream. On the other hand, NGC 4725 was identified as a double-barred galaxy by Erwin (2004), but de Lorenzo-Cáceres et al. (2013) suggested that, at present, the inner bar is playing a minor role in promoting major morphological changes. Chiaraluce et al. (2019) attempted to measure its nuclear radio emission, including this object among a sample of low-luminosity Seyfert galaxies (Véron-Cetty & Véron, 2006). However, their Very Large Array (VLA) observations produced an unresolved image and a peculiar radio spectrum, which merits further investigation to derive meaningful information from this source.
@@
Regarding NGC 5236, this galaxy has a grand-design barred spiral morphology and hosts a nuclear starburst ring (Buta & Crocker, 1993; Calzetti et al., 2004; Knapen et al., 2010) leading to a high total SFR of 4.2 yr-1 (Leroy et al., 2021). Furthermore, Della Bruna et al. (2022) observed that both the stars and gas kinematics of this galaxy trace regular rotation along the main star-forming disc as well as a fast-rotating circumnuclear disc of ( 700 pc) in diameter, likely originating from secular processes driven by the galactic bar (as already reported by Gadotti et al. (2020)). Della Bruna et al. (2022) also detected an extended ionised gas region, which could be attributed to a bar-driven inflow of shocked gas, among other factors.
@@ 2.4 Environmental properties
The surrounding environment of galaxies is believed to play a crucial role in their evolutionary path (Dressler, 1980; Poggianti et al., 1999). According to many observational works, late-type galaxies (spiral, blue, and star forming) are more likely to be found in lower density environments than early-type ones (elliptical, red, and quenched), which are more numerous in groups, clusters, and interacting systems (e.g. Casasola et al., 2004; Blanton et al., 2005; Peng et al., 2010; Boselli & Gavazzi, 2014). Indeed, there are several external processes that are commonly identified as responsible for this phenomenology and that can alter the effects of internal and/or secular processes, such as the action of a stellar bar. Among them, we can mention: ram-pressure stripping (Gunn & Gott, 1972; Yun et al., 2019), mergers (Mihos & Hernquist, 1994), harassment (high-speed encounters between galaxies; Moore et al., 1996), and strangulation (cutoff of gas accretion; Kawata & Mulchaey, 2008). Below, we present the environment in which each of our barred galaxies is located; we investigated this aspect in order to search for any possible influence of environment on star formation.
@@
NGC 1566 is the brightest member of the Dorado group (Agüero et al., 2004; Kilborn et al., 2005) and belongs to a pair-like structure where its companion is the dwarf elliptical NGC 1581 (Rampazzo et al., 2022). The work of Elagali et al. (2019) found that NGC 1566 has an asymmetric and mildly warped HI disc. To explain this feature, these authors excluded the scenario of past interactions with other galaxies in favour of the possibility of having experienced ram-pressure stripping from the intergalactic medium (IGM), at least in its outskirts; however, the high atomic gas fraction that characterises this galaxy hints at a subtle ram-pressure stripping effect, as expected in galaxy groups.
@@
We obtained information on the environmental properties of NGC 3351 and NGC 5236 from the catalogue produced by Karachentsev et al. (2013). This latter contains information about 869 galaxies within 11 Mpc and includes three estimates of the ‘tidal index’, which quantifies the local density environment. In particular, we looked at the tidal index , which is determined by the five most important neighbours. Positive values of this parameter indicate group membership, while negative values correspond to either isolated galaxies or galaxies in the outskirts of a group. NGC 3351 has a relatively high value of compared to the distribution of galaxies in the Karachentsev et al. (2013), meaning that it is in a mildly dense environment. On the other hand, NGC 5236 has and so is considered a ‘field’ galaxy. In support of this definition, we mention that NGC 5236 is present in the catalogue of isolated galaxies of Bettoni et al. (2003).
@@
For what concerns NGC 3953, Kourkchi & Tully (2017) identified it as part of a group, whose main member is the barred spiral NGC 3992. According to Vollmer et al. (2012), NGC 4579 is one of the brightest galaxies in the Virgo Cluster. It is an anemic galaxy and has a relatively small amount of neutral hydrogen located inside its disc and concentrated in clumps (Cayatte et al., 1990). This deficiency of gas content is thought to be a consequence of interactions with the intracluster medium of the Virgo Cluster, suggesting that ram-pressure stripping may have occurred in the galaxy’s outskirts (Koopmann & Kenney, 2004).
@@
NGC 4725 is the brightest member of the Coma I aggregation (Gregory & Thompson, 1977) and seems to be interacting with the neighbouring spiral galaxy NGC 4747 (Haynes, 1979). Gravitational interactions between these two galaxies likely resulted in a gas-rich tidal tail that extends about ( 30 kpc) to the northeast of NGC 4747 and appears to contain two distinguishable star clusters (Lee-Waddell et al., 2016, 2018). On the other hand, only a relatively small distortion in the HI distribution is observed in NGC 4725.
@@ 2.5 Multi-wavelength observations
As briefly mentioned above, DustPedia provides multi-wavelength observations from UV to microwave. Regarding both Spectral and Photometric Imaging REceiver (SPIRE; Griffin et al., 2010) and Photodetector Array Camera and Spectrometer (PACS; Poglitsch et al., 2010) data, the Herschel Science Archive (HSA555https://archives.esac.esa.int/hsa/whsa/) was queried to find all Herschel photometric observations covering each DustPedia galaxy. The ancillary imaging data were gathered from a further seven facilities that observed large numbers of nearby galaxies: the GALaxy Evolution eXplorer (GALEX; Morrissey et al., 2007), the Sloan Digital Sky Survey (SDSS; York et al., 2000; Eisenstein et al., 2011), the Digitized Sky Survey (DSS), the 2 Micron All-Sky Survey (2MASS Skrutskie et al., 2006), the Wide-field Infrared Survey Explorer (WISE Wright et al., 2010), the Spitzer Space Telescope (Werner et al., 2004), and Planck (Planck Collaboration I, 2011). More information on the different data sets available can be found in Enia et al. (2020) and references therein.
Thanks to the data derived from these multi-wavelength observations collected in the DustPedia archive, it is possible to exploit the power of complete multi-band photometric coverage, and especially its ability to provide a measure of the comprehensive SFR in galaxies, directly taking into account both the optical and dust-obscured components. For what concerns our work, we use observations from GALEX, SDSS, 2MASS, WISE, Spitzer, and Herschel, going from the FUV (1516 Å) to the FIR (350 m). However, it has to be noted that not all of these data sets are accessible for the whole sample. Indeed, the entire set of 23 photometric bands is only available for NGC 3351, NGC 4579, and NGC 4725. NGC 3953 is observed in 20 bands; GALEX filters (FUV, NUV) and PACS 70m photometry are not available for this object. Instead, 18 bands are accessible for NGC 1566 and NGC 5236, which lack optical observations (the five Sloan optical filters).
Figure 1 shows the selected galaxies as they are observed in six different bands: FUV with GALEX, g-band with SDSS, J-band with 2MASS, 4.5 m with Spitzer, 12 m with the WISE, and 250 m with Herschel. An elongated morphological structure is clearly visible in all the galaxies, both in the SDSS optical bands (when available) and in the NIR ones (2MASS and Spitzer).
@@ 3 Data analysis
The primary goal of our data analysis is to obtain the spatially resolved distributions of the two galaxy physical properties: stellar mass and SFR. In this respect, we carried out a SED fitting procedure on subkpc scales using the publicly available magphys code (da Cunha et al., 2008). The methodology that we adopt to perform the data analysis has been developed and already used in previous works for a sample of local star forming galaxies with a grand-design spiral structure (Enia et al., 2020; Morselli et al., 2020). Our aim is to ensure a fair comparison with their results for unbarred galaxies, as they have consistent global properties with respect to ours.
In Enia et al. (2020), the procedure is summarised in three main steps: (i) degradation of the resolution of each image to the worst point spread function (PSF) that is available among all the photometric bands (i.e. that of the m image taken with the SPIRE instrument, characterised by a PSF FWHM of ); (ii) the setting up of a grid of square cells of fixed size from which we measure the flux at different wavelengths; and (iii) derivation of the physical properties for individual apertures by performing SED fitting to the available photometric data. We analyse every selected source considering cells of (corresponding to physical scales varying from to , as reported in Table 1), which is the pixel scale of SPIRE m maps. This choice is justified by the fact that we need a good balance between resolution and the number of available photometric bands in order to properly sample the SED —including the dust emission curve— and to avoid an excessively low resolution, which would prevent us from investigating the bar region. Specifically, it is necessary to sample the bar radius of each galaxy through at least three resolution elements (cells of each) in order to obtain accurate information about this morphological structure. At this point, we highlight the fact that, throughout the paper, we refer to the approximately annular region around the very central nucleus that extends up to the radius of the bar as the ‘region hosting the bar’.
@@ 3.1 Pre-processing and PSF degradation
The preliminary steps that are performed on each photometric observation are those related to the preprocessing and PSF homogenisation phases, which start with the removal of foreground stars by masking contaminating sources through an automated procedure embedded in CAAPR. Subsequently, a background estimation and subtraction are performed on each of the star-subtracted maps following the guidelines provided by Clark et al. (2018). Finally, the resolution of all the stars and background-subtracted maps is degraded to the PSF of SPIRE 350, exploiting the kernels of Aniano et al. (2011).